# Privacy-Aware In-Context Learning for Large Language Models

### v3

@@ Appendix A Proof of Theorem 1
Appendix B Hyperparameter tuning
@@
TABLE V: Values for hyperparameters explored in this work.
Appendix C Empirical privacy evaluation by membership inference attacks
@@
TABLE VI: Empirical privacy evaluation of our method for 1-shot ICL by MIA on Gemma 1.1 2B IT model.
While differential privacy offers theoretical privacy guarantees, empirical evaluation is also crucial [64]. To assess the practical privacy of our DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) [66], a practical method for measuring real-world privacy leakage.
We follow [20] to instantiate MIA in the in-context learning (ICL) setting, where the goal is to detect whether a data point was included in the LLM prompt. As expected, using true private samples leads to successful attacks. To evaluate our DP few-shot generation, we split the DBPedia dataset into member and non-member sets. Using member data, we generate 1-shot demonstrations with our DP algorithm for $\epsilon=4$ on 5 runs. For MIA, we query 50 member and 50 non-member samples, repeating for 100 trials to compute the average AUC. For the non-private baseline, we follow the same setup using actual member samples in the prompt.
Table VI shows the MIA results. Consistent with [20], using actual private samples yields high AUC (94.20 for $\epsilon=\infty$), indicating successful attacks. In contrast, our DP approach significantly reduces the AUC, demonstrating improved membership privacy.
Appendix D Effect of model size/architecture
$\varepsilon=1$ Model Acc. (%)
google/gemma3-1b-it $67.0$
google/gemma2-2b-it $84.9$
meta-llama/Llama-3.2-1B $53.9$
meta-llama/Llama-3.2-3B $84.2$
meta-llama/Meta-Llama3-8B $89.9$
TABLE VII: DBPedia classification accuracy for various LLMs. All evaluations use the same ICL setup described in §Experimental Setup.
Appendix E Generated Synthetic Examples
@@
Appendix F Public and Private Prompts

### v4

@@ 9 $x_{t}\sim\operatorname{softmax}(\hat{z}/\tau)$
Experimental Setup
Datasets:
For evaluation, we follow the prior ICL work (zhao2021calibrate) and use the following setup. For classification tasks, we use three datasets: the 4-way news classification dataset AGNews (zhang2015character), the 6-way question classification dataset TREC (zhang2015character), and the 14-way topic classification DBPedia (voorhees2000building). We also evaluate on two information extraction tasks, namely MIT-G, and MIT-D (liu2012conversational). These are both slot-filling datasets with movie genre (MIT-G) and director name (MIT-D) as the slots to be filled. An illustrative example is shown in Figure 2.
Since our method requires access to the full token probability distribution at each decoding step, we use the instruction-tuned (IT) variant of Gemma 1.1 (gemmateam2024gemma), a decoder-only LLM with two billion parameters, aligning with prior work by amin2024private. However, this choice is made purely for experimental benchmarking. Our algorithm is model-agnostic and can be applied to any decoder-style LLM, provided access to full token-level logit distribution is available. We provide additional details for chosen hyper-parameters in the Appendix.
Results
Table 1 presents our in‐context learning results on the five benchmark tasks. Our primary choice of LLM is Gemma-1.12B IT which provides access to model logits as opposed to closed models such as “GPT3-baggage” used by some previous works (tang2024privacypreserving). We evaluate and compare the utility of our model with varying level of privacy budget ($\varepsilon=0,\infty,1$). The first block with $\varepsilon=0$ has the highest amount of privacy shows zero-shot performance (no ICL) is uniformly poor across tasks (e.g., 24.8% on AGNews).
The second block with $\varepsilon=\infty$ examines when the least amount of noise is added in our DP framework. We observe that we perform better than the current SOTA amin2024private’s baseline on Gemma-1.12B in four tasks, namely DBPedia, TREC, MIT-G and MIT-D, while lagging behind by only $3.3\%$ on AGNews. For most datasets: DBPedia, TREC, MIT-G and MIT-D, we also outperform tang2024privacypreserving’s baseline with GPT3 Babbage that uses only top-100 logprobs.
The third block with $\varepsilon=1$ imposes stronger privacy constraints on synthetic data. We observe that we outperform both baselines by tang2024privacypreserving, and  amin2024private with Gemma-1.12B on three out of five tasks: AGNews, TREC, MIT-G. Similarly, for AGNews, TREC, MIT-G and MIT-D datasets, we outperform tang2024privacypreserving’s baseline with GPT3 Babbage that uses only top-100 logprobs.
Effects of $k$-Shot and Privacy Budget Settings
$\epsilon$ 1-shot 2-shot 4-shot 8-shot 12-shot
1 81.66 82.46 76.80 74.82 74.46
4 79.90 82.62 79.84 76.04 68.52
8 77.00 83.34 78.62 74.90 68.56
$\infty$ 80.86 80.48 80.94 81.44 75.48
Table 2: ICL accuracy under varying privacy budgets with different $k$-shots on DBPedia.
$\epsilon$ 1-shot 2-shot 4-shot 8-shot 12-shot
1 68.32 60.60 63.00 64.40 65.84
4 68.76 59.72 62.48 63.08 63.28
8 68.96 62.00 63.20 67.08 62.32
$\infty$ 62.40 54.04 61.92 67.16 70.12
Table 3: ICL accuracy under varying privacy budgets with different $k$-shots on TREC.
In Table 2, results on DBPedia show that performance generally improves with larger values of $\epsilon$, indicating that relaxing privacy constraints allows for more effective use of private data in generating synthetic demonstrations. However, as the number of shots increases, the performance can sometimes degrade slightly possible due to the lack of targeted private signal which can also be the training artifacts of LLM training. Finally, the best performance for 4-shot (80.94), 8-shot (81.44), and 12-shot (75.48) appears in the $\epsilon=\infty$ (non-private) case, highlighting the utility-privacy trade-off inherent to DP algorithms.
Table 3 presents the corresponding analysis on the TREC dataset, where the pattern across privacy levels is more nuanced. For example, $\epsilon=8$ outperforms all other settings for it should be 1-shot, 2-shot, and 4-shot configurations, indicating that a moderately relaxed privacy guarantee yields significant utility benefits in certain tasks.
Structured Data Generation
We evaluate how well our approach preserves privacy while generating syntactic structures from sensitive data. In structured generation tasks, many tokens are essential for maintaining the correct output structure. To assess this, we experiment on the WikiMoviesJSON generation task, using preprocessing and evaluation setups described in (amin2024private). We assess performance using two metrics: (1) the percentage of outputs that are syntactically well-formed JSON parses, and (2) the percentage of outputs that pass basic schema validation. As shown in Table 4, our method achieves high-quality and schema-compliant JSON generation even under a strict privacy budget of $\epsilon=1$, demonstrating the effectiveness of our approach for privacy-preserving structured text synthesis.
$\varepsilon$ Method $\tau$ Parses (%) Validates (%) #\#raw
1 amin2024private 2 $80.6_{1.3}$ $74.2_{1.9}$ $94.3_{1.2}$
2.5 $4.9_{1.1}$ $1.5_{0.1}$ $138.0_{7.5}$
Ours 1.13 $84.2_{4.08}$ $81.1_{8.1}$ $10.3_{1.15}$
Table 4: Results for generating JSON records from WikiMoviesJSON and report gains in structure preservation and validations. We report mean and standard deviation over 3 runs of dataset generation. Here, $\tau$ refers to the sampling temperature, and #\#raw refers to the number of raw samples produced before parsing and validation checks.
Effect of Model Size and Architecture
Model Acc. (%)
google/gemma3-1b-it $67.0$
google/gemma2-2b-it $84.9$
meta-llama/Llama-3.2-1B $53.9$
meta-llama/Llama-3.2-3B $84.2$
meta-llama/Meta-Llama3-8B $89.9$
Table 5: DBPedia classification accuracy for various LLMs. All evaluations use the same ICL setup described in §Experimental Setup and the privacy budget is set to $\varepsilon=1$.
Privacy Attacks
While differential privacy offers theoretical privacy guarantees, empirical validation is essential (blanco2022critical). We conduct personally identifiable information (PII) extraction attack on the Enron email dataset following the experimental setup described in (zeng2024good). Specifically, we first construct a private dataset of text which contain email addresses in the body. Using Algorithm 1, we construct a private subset $S$ and generate $T=15$ synthetic tokens using the prompt template as “Extract only the email address from the above text.”. The goal of this attack is to test whether our approach prevents release of private email addresses in the generated text. We evaluate under various privacy budgets, namely $\epsilon=[1,4,8]$. Across all tested privacy levels, we observe zerozero sensitive email address in the generated responses demonstrating the effectiveness of our approach in preventing leakage of PII and demonstrate its strong practical privacy-preserving capabilities.
Furthermore, to assess the practical privacy of DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) (ShokriSSS17) following (duan2023flocks). When actual private samples are used in prompts, attacks succeed with a high AUC (94.20 for $\epsilon=\infty$). In contrast, our DP approach significantly reduces AUC ($51.78$ for $\epsilon=4$), confirming improved membership privacy. See Appendix Table MIA for details.
Discussion
Traditional differentially private generation methods, such as those employed by amin2024private, rely on metric-based mechanisms like the Sparse Vector Technique (SVT), which requires threshold computations under additive noise. These approaches typically employ distributional distance metrics such as $\ell_{1}$ to score or select candidate tokens. However, these metrics are computed over normalized probability distributions (i.e., post-softmax), where even semantically similar tokens with slightly different probabilities or indices can yield high $\ell_{1}$ distance values. This issue is especially pronounced in high-dimensional output spaces of language models, where such metrics treat tokens as orthogonal dimensions, ignore semantic similarity, and are insensitive to token ranking that are crucial for meaningful text generation.
Moreover, approach by amin2024private continues generating additional sentences beyond termination, often yielding low-quality synthetic examples. Our procedure strictly terminates when <eos> token is encountered. This guarantees that all outputs are semantically coherent and suitable for downstream tasks. Qualitative examples supporting this claim are provided in the Appendix.
Conclusion
As access to foundation models grows, the resources required for training these models have become expensive; hence, private prediction could emerge as a compelling alternative to private fine-tuning. In this work, we show that private prediction can generate synthetic text while following the standard differential-privacy guarantees. This privately generated corpus substantially boosts performance in many-shot in-context learning. Moreover, introducing a mechanism for sampling tokens from public models and blending them with private tokens enhances the utility of prediction tasks without compromising the privacy accounting.
Limitations
Acknowledgment
Appendix A Hyperparameter tuning
@@
Table 6: Values for hyperparameters explored in this work.
Appendix B Empirical privacy evaluation by membership inference attacks
@@
Table 7: Empirical privacy evaluation of our method for 1-shot ICL by MIA on Gemma 1.1 2B IT model.
While differential privacy offers theoretical privacy guarantees, empirical evaluation is also crucial (blanco2022critical). To assess the practical privacy of our DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) (ShokriSSS17), a practical method for measuring real-world privacy leakage.
We follow (duan2023flocks) to instantiate MIA in the in-context learning (ICL) setting, where the goal is to detect whether a data point was included in the LLM prompt. As expected, using true private samples leads to successful attacks. To evaluate our DP few-shot generation, we split the DBPedia dataset into member and non-member sets. Using member data, we generate 1-shot demonstrations with our DP algorithm for $\epsilon=4$ on 5 runs. For MIA, we query 50 member and 50 non-member samples, repeating for 100 trials to compute the average AUC. For the non-private baseline, we follow the same setup using actual member samples in the prompt.
Table 7 shows the MIA results. Consistent with (duan2023flocks), using actual private samples yields high AUC (94.20 for $\epsilon=\infty$), indicating successful attacks. In contrast, our DP approach significantly reduces the AUC, demonstrating improved membership privacy.
Appendix C Generated Synthetic Examples
@@
Appendix D Public and Private Prompts